WARNING: This web is not used anymore. Please use PDBService.MinuteS15May08 instead!
 

Minutes 15 May 08

  • Phone:
    • Gridka: Doris
    • RAL: Gordon
    • NDGF: Olli
    • BNL: Carlos, John
    • TRIUMF: Andrew
    • SARA: Alexander
    • Taiwan: Dave
    • PIC: Luis
    • CMS: Pietro

  • CERN:
    • 3D: Dirk, Maria, Eva
    • ATLAS: Gancho, Florbela

  • Site status:
    • RAL: testing new hw which has arrived
    • NDGF: nothing to report
    • Gridka:
      • Replication problem from LHCb to Gridka, propagation is aborted with connection lost contact error. SR opened with Oracle.
      • Some problems with FTS cluster. Problems installing FTS monitoring, Doris has contacted FTS people.
      • BNL: nothing to report
      • Taiwan: deadlock problem with asm instance. Dirk suggested to contact to Castor support because the problem was observed (castor deployment list).
      • Triumf: nothing to report
      • PIC: PIC new setup for LHCb. ATLAS migration proposed to first week of June, cluster database is being prepared. LFC for ATLAS will be migrated 2nd week of June.
      • SARA: apply process aborted two weeks ago: no data found. The accounts were not locked at the destination. Rows being updated at the same time at source and destination. Fixed.

  • Experiment status:
    • ATLAS:
      • Monitoring of destination databases: Florbela will re-send the link (include in the agenda)
      • PVSS schema instantiation test between INTR and INT8R using transportable tablespaces. Production end of June.
      • AMI replication from Lyon to CERN: use INTR for the test environment. Contact them to register to the 3d list + configuration check list for source database + streams administrator user
      • Muon sites: 3 sites to be replicated to CERN. First tests using INTR.
      • Problem between online and offline. Primary keys were missing for a set of tables and one table had all the rows duplicated. Apply process was enabled to find the appropriate row. Duplicated rows were cleaned and new copy of tables imported at the destination.
    • CMS:
      • Online to Offline replication being tested.

  • Carlos (BNL): Question on applying patch 10.2.0.4 + cpu patch. We have to wait until the CCRC May is finished.

  • Meeting in 3 weeks.

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2008-05-16 - EvaDafonte
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback